OpenStudy (anonymous):

Suppose that the graph of f contains the point (6,4). Find a point that must be on the graph of g(x)=3+f(5x).

OpenStudy (anonymous):

similar to the last one if \(5x=6\) then what is \(x\) ?

OpenStudy (anonymous):

x=1.2

OpenStudy (anonymous):

ok, and if \(x=1.2\) then what is \(3+f(5\times 1.2)\)?

OpenStudy (anonymous):

how do you get to 5x=6?

OpenStudy (anonymous):

you know what \(f(6)\) is right?

OpenStudy (anonymous):

and that is all you know about \(f\)

OpenStudy (anonymous):

so if you have any hope of knowing what \(3+f(5x)\) is, you can only know it if \(5x=6\)

OpenStudy (anonymous):

okay i understand that part now

OpenStudy (anonymous):

when does the 4 come in?

OpenStudy (anonymous):

you know that \(f(6)=4\)

OpenStudy (anonymous):

so if \(x=1.2\) then \(3+f(5x)=3+f(5\times 1.2)=3+f(6)=3+4=7\)

OpenStudy (anonymous):

meaning \((1.2,7)\) is on the graph of \(g\)

OpenStudy (anonymous):

okay thank you im gonna try to next one on my own to see if i can do it

OpenStudy (anonymous):

i have another problem where f contains point 5,2 and g(x)=7(f(x)+3) i found x=5 but i dont know what y is i thought it was 14 but my online homework says its wrong